Use inductive reasoning to predict the next two terms in the pattern. Justify your prediction


1, 8, 27, 64, 125...

Answer by Theo(13342)

1^3 = 1
2^3 = 8
3^3 = 27
4^3 = 64
5^3 = 125
6^3 = 216

The next term is 216.
Every term is equal to the next number in the sequence cubed.